OpenStudy (anonymous):

Write the equation of the given line in standard form. Coordinates are (1/2, 3) and (-4,-1) A. -8x + 9y = 23 B. -8x + 9y = -23 C. -8x + 7y = 25 D. -9x + 8y = -23 The answer is A

OpenStudy (anonymous):

If you want to simply solve this problem, just plug in 1 of the x and y cord pairs into the equations until you find one that works

OpenStudy (anonymous):

If you want to do it the hard way... 1st you have to rewrite all the equations so that y is by itself on 1 side (y=mx+b) Of course...that's way too much work on an exam

OpenStudy (anonymous):

i got C... am i right?

OpenStudy (anonymous):

B.. sorry hehe

OpenStudy (anonymous):

uhh...lemmme check, haven't done it yet :P

OpenStudy (anonymous):

oh okay haha sorry, no rush

OpenStudy (anonymous):

It's actually A -8 x -4 + 9 x -1 = 23 32 - 9 = 23